OpenStudy (anonymous):

a certain car is capable of accelerating at a rate of 0.60 m/s^2. How long does it take for this car to go from a speed of 55mi/h to a speed of 60 mi/h?

OpenStudy (anonymous):

use conversion a=0.6m/s^2 X 3.28ft/m=1.968 ft/s^2 V0=55X5280/3600 =80.67 FT/SEC V1=50X5280/3600=88 FT/SEC now use the formula at=v1-v0 t=(88-80.67)/1.968 t=3.72 sec .....ans